OpenStudy (anonymous):

Find all the zeros of the function f(x)= (x+5)(x-8)^2

OpenStudy (anonymous):

the (only) values for which f(x) is zero is when (x+5) is zero or when (x-8) is zero; since zero time anything is zero. equations x + 5 = 0 and x - 8 = 0 will find both zeros

OpenStudy (anonymous):

Since (x+5) and (x-8) are the two factors of this polynomial therefore (x+5) = 0 and (x-8) =0 i.e. x=-5 and x=8 are the two zeroes of the given polynomial.